# Inspection-Guided Randomization: A Flexible and Transparent Restricted Randomization Framework for Better Experimental Design

### v1

@@ 4. Evaluation & Adaptation
    Desiderata trade-offs: When there are multiple competing design desiderata, it may not be immediately clear whether and how to best construct the fitness function to weigh desiderata trade-offs. Visualizing the space of all candidate allocations by plotting the score outputs of different inspection metrics on separate axes can assist in understanding the nature of desiderata trade-offs. Then, visualizing the space of accepted allocations can assist in understanding whether the fitness function navigates such trade-offs as desired. If necessary, the analyst can make informed adjustments to the fitness function by modifying the aggregator function $\Theta$ that combines the inspection metric scores. In Section 4.2, we provide an example of how adjusting the prioritization weights for the individual inspection metrics can be used to navigate trade-offs between achieving covariate balance and minimizing interference between treatment arms.
@@
    Over-restriction: We consider two types of over-restriction. In the first type of over-restriction, the combination of fitness function and restriction rule may result in too few accepted allocations to obtain p-values with the desired level of granularity (see Section 3.4 for details on deriving p-values from randomization tests). When this occurs, the analyst may need to modify the restriction rule to be less aggressive or choose to apply a restriction rule that always selects the top $m$ allocations. Too few accepted allocations could also be a consequence of using an enumeration strategy that does not “hit” the correct regions of the space of all possible allocations.This issue can be addressed by increasing the size of the enumerated pool or by using stochastic optimization routines (see Section 3.3). In the second type of over-restriction, the set of accepted allocations may be too similar to one another, such that multiple units or groups of units are assigned to the same treatment arm across most accepted allocations. If we rely on an asymptotic p-value in testing the null hypothesis, then this kind of over-restriction can result in departures from the nominal Type I error rate (Moulton,, 2004). An exact p-value obtained through randomization inference circumvents this issue, but highly correlated allocations can also lead to increased mean-squared error and reduced power, regardless of whether randomization inference is used (Nordin and Schultzberg,, 2022; Krieger et al.,, 2020). Relaxing the restriction rule $r$ to permit additional allocations is one way to alleviate over-restriction issues.
5. Pre-registration
When the research team is satisfied with the choice of fitness function and restriction rule, they lock in the assignment mechanism $P_{IGR}$ and pre-register the fitness function $f$, the restriction rule $r$, and the pool of candidate allocations $\mathcal{Z}_{pool}$ in the trial protocol. If the number of candidate allocations is large, it may be more practical to pre-register the code, including the setting of a particular random seed, used to generate the pool of candidate allocations.
6. Randomization
Finally, the observed allocation $\mathbf{z}^{obs}\sim P_{IGR}$ is drawn and deployed.
3.3 IGR enhancements
3.3.1 Enhancement 1: Using genetic algorithms for efficient enumeration
Since $M<<k^{N}$, it is possible that many desirable, high-scoring allocations are overlooked and not enumerated as part of $\mathcal{Z}_{pool}$. Instead of using simple enumeration strategies, we may instead want to efficiently search through the space of all $k^{N}$ possible allocations to construct a more optimal $\mathcal{Z}_{pool}^{(*)}$. One approach to construct $\mathcal{Z}^{(*)}_{pool}$ is to use genetic algorithms (Mitchell,, 1998). The initial $\mathcal{Z}_{pool}$ is viewed as a population of treatment allocations. These allocations “mate” in pairs to produce a new generation of allocations, where novel allocations are introduced through “mutation” (single entries of an allocation are changed) and “crossover” (segments of two parent allocations are swapped). The individuals in this generation are scored using the specified fitness function $f$, and those with the best scores survive and breed to produce the next generation. We can repeat this process over $T$ generations to obtain $\mathcal{Z}_{pool}^{(*)}$. Henceforth, we use “IGRg” to refer to IGR enhanced by the use of genetic algorithms in the enumeration step.
3.3.2 Enhancement 2: Adding mirror allocations
@@
Assuming no interference, the difference-in-means estimate is unbiased for the average treatment effect when the assignment mechanism satisfies the mirror property. For a binary intervention, the mirror property is satisfied if $P(Z_{i}=0)=P(Z_{i}=1)$ for all $i=1,\dots,N$. Suppose we enumerated all possible allocations in the Enumeration step. If we specify fitness function $f$ and restriction rule $r$ to be symmetric, such that $(r\circ f)(\mathbf{X},\mathbf{z}))=(r\circ f)(\mathbf{X},\mathbf{1}-\mathbf{z}))$, the mirror property will be satisfied by $P_{IGR}$. However, because we may not be able to enumerate all possible allocations, satisfaction of the mirror property is not guaranteed. Provided $f$ and $r$ are symmetric, we can update $P_{IGR}$ by ensuring that, for every allocation that is accepted, its mirror allocation is also accepted. Let $\tilde{\mathcal{Z}}_{pool}=\{1-\mathbf{z}\ |\mathbf{z}\in\mathcal{Z}_{pool}\}$ denote the set of mirror allocations for each allocation in $\mathcal{Z}_{pool}$. The assignment mechanism $\tilde{P}_{IGR}$ satisfies the mirror property, where
@@
We can also naturally extend this mirroring to a multi-arm experiment in which one arm is designated the control arm and the remaining are treated.
In some experiments, the design desiderata may not be symmetric. For example, in an experiment with interference, the interference itself can create bias that is not resolved by satisfaction of the mirror property. Thus, choosing an “asymmetric” $f,r$ pair that does not satisfy the mirror property may sometimes help with reducing interference bias. However, asymmetry can also result in a violation of marginal independence between the treatment assignment variable and the potential outcomes. If $f,r$ is not perfectly symmetric, but the difference between $(r\circ f)(\mathbf{X},\mathbf{z})$ and $(r\circ f)(\mathbf{X},\mathbf{1}-\mathbf{z})$ is small for all $\mathbf{z}$, then the analyst may consider still incorporating the mirror allocations $\mathcal{\tilde{Z}}_{pool}$. We note that, while marginal independence may no longer hold under asymmetry, the treatment assignment variable will always be conditionally independent of the potential outcomes, where the conditioning set consists of the covariates used to calculate the fitness score. Thus, if not all mirrors are included, then a form of covariate adjustment can be performed during the analysis phase.
3.4 Analyzing IGR experiments: Randomization inference

### v2

@@ 4.2.4 Evaluation & Adaptation
    Desiderata trade-offs: When there are multiple competing design desiderata, it may not be immediately clear whether and how to best construct the fitness function to weigh desiderata trade-offs. Visualizing the space of all candidate allocations by plotting the score outputs of different inspection metrics on separate axes can assist in understanding the nature of desiderata trade-offs. Then, visualizing the space of accepted allocations can assist in understanding whether the fitness function navigates such trade-offs as desired. If necessary, the analyst can make informed adjustments to the fitness function by modifying the aggregator function that combines the inspection metric scores.
@@
    Over-restriction: We consider two types of over-restriction. In the first type of over-restriction, the combination of fitness function and restriction rule may result in too few accepted allocations to obtain p-values with the desired level of granularity (see Section 5 for details on deriving p-values from randomization tests). When this occurs, the analyst may need to modify the restriction rule to be less aggressive or choose to apply a restriction rule that always selects the top $m$ allocations. Too few accepted allocations could also be a consequence of using an enumeration strategy that does not “hit” the correct regions of the space of all possible allocations. This issue can be addressed by increasing the size of the enumerated pool or by using stochastic optimization routines. In the second type of over-restriction, the set of accepted allocations may be too similar to one another, such that multiple units or groups of units are assigned to the same treatment arm across most accepted allocations. If we rely on an asymptotic p-value in testing the null hypothesis, then this kind of over-restriction can result in departures from the nominal Type I error rate (Moulton,, 2004). An exact p-value obtained through randomization inference circumvents this issue, but highly correlated allocations can also lead to increased mean-squared error and reduced power, regardless of whether randomization inference is used (Nordin and Schultzberg,, 2022; Krieger et al.,, 2020). Relaxing the restriction rule $r$ to permit additional allocations is one way to alleviate over-restriction issues.
{leftrule}
Evaluation & Adaption in a Simulated Experiment
We use visualizations to assist in evaluating our choice of fitness function and restriction rule. Additional toy examples and a more detailed discussion on the choice of visualizations can be found in Appendix D. Below, we show an example iterative process of evaluation and adaptation for our simulated experiment.
Iteration 1: To assess discriminatory power, we plot a histogram of the scores for the candidate allocations and check that the histogram has sufficient spread (Figure 1, left). A “spiky” histogram where all fitness scores are concentrated around a few values is indicative of a fitness function that has low discriminatory power.
To examine desiderata trade-offs in the pool of candidate allocations, we plot a two-dimensional histogram with the balance metric on the x-axis and the interference metric on the y-axis (Figure 1, middle). By looking at this heatmap, we can check whether allocations with better values for one metric tend to have worse values in the other metric. To examine desiderata trade-offs in the accepted set of candidate allocations, we make a scatterplot with the balance metric on the x-axis and the interference metric on the y-axis. Each point in the scatterplot corresponds to a single accepted allocation. By comparing the positioning of the accepted allocations to those of a benchmark randomization strategy, we can evaluate whether our restricted randomization scheme is navigating desiderata trade-offs as desired.
To check for over-restriction, we plot a histogram of the approximate pairwise assignment correlation (Figure 1, right). Specifically, for each pair of units, $i,j$, we compute the fraction of candidate allocations where $i$ and $j$ are assigned to the same arm. When there is no pairwise assignment correlation, then each pair $i,j$ should appear in the same arm in half of the candidate allocations.
Figure 1: Visual checks for discriminatory power, desiderata trade-offs, and over-restriction (via approximate pairwise assignment correlation) under fitness function $f^{(\texttt{I+M})}$.
{leftrule}
Iteration 2: Based on the score histogram in Figure 1, it appears that the fitness function provides sufficient differentiation between allocations. Based on the desiderata trade-off plots, however, we can see that the interference metric is extremely coarse and takes on only a handful of distinct values. At the smallest value of the interference metrics, fine-grained fitness score differentiation is driven primarily by the balance metric. This may be appropriate if we are only interested in filtering out the most interference-prone allocations. If, on the other hand, we want to be more selective against interference even in the low-interference regimes, then a different interference metric is necessary.
Suppose we are able to collect more detailed knowledge of a network of interactions between individuals. For instance, we can ask participants to list the friends with whom they communicate regularly and build a network based on each participant’s list of contacts (Cai et al.,, 2015; Kim et al.,, 2015). In the network interference literature, it is common to assume that an individual is “exposed” through the treatments of their neighbors in the network via a neighborhood exposure model (Aronow and Samii,, 2017; Manski,, 2013). We therefore propose a new interference metric that counts the fraction of control individuals exposed to treatment through their network neighbors,
$m^{(\texttt{FracExpo})}(\mathbf{z},\mathbf{A})=\frac{\sum_{i=1}^{N}\mathbf{1}\{\sum_{j\in\mathcal{N}(i)}z_{j}>q\cdot|\mathcal{N}(i)|\}\cdot(1-z_{i})}{\sum_{i=1}^{N}(1-z_{i})}$ (4)
[…]
$\begin{split}f^{(\texttt{F+M})}(\mathbf{z},\mathbf{X})&=0.5\cdot m^{(\texttt{FracExpo})}(\mathbf{z},\mathbf{X})+0.5\cdot m^{(\texttt{Mahalanobis})}(\mathbf{z},\mathbf{X})\end{split}$ (5)
We re-make each of the plots from iteration 1 (Figure 2). From the desiderata trade-off plots we see that, as desired, the new interference metric based on network exposure is more granular than the metric based on distance. We also see that the new overall fitness function has adequate discriminatory power and has similar approximate pairwise assignment correlation to complete randomization.
Figure 2: Visual checks for discriminatory power, desiderata trade-offs, and over-restriction (via approximate pairwise assignment correlation) under fitness function $f^{(\texttt{F+M})}$.
{leftrule}
Iteration 3: We decide to use the interference metric based on network exposure. In previous iterations, the interference metric and balance metric were weighted equally. Now, we want to examine an alternative weighting scheme where more weight is applied to the interference metric. We therefore propose the fitness function,
$\begin{split}f^{(\texttt{F+M})}(\mathbf{z},\mathbf{X})&=0.75\cdot m^{(\texttt{FracExpo})}(\mathbf{z},\mathbf{X})+0.25\cdot m^{(\texttt{Mahalanobis})}(\mathbf{z},\mathbf{X})\end{split}$ (6)
Again, we construct plots to check for discriminatory power, desiderata trade-offs, and over-restriction (Figure 3). From the desiderata trade-offs scatterplot, we can see that a fitness function that upweights the interference metric accepts allocations with fewer exposed control units, at the cost of increased imbalance on covariates. The pairwise assignment correlation remains similar to complete randomization.
Figure 3: Visual checks for discriminatory power, desiderata trade-offs, and over-restriction (via approximate pairwise assignment correlation) under fitness function $f^{(\texttt{F+M})}$ with an upweighted interference metric.
{leftrule}
Final decision: Ultimately, we proceed with the fitness function from Iteration 3 and the restriction rule used across iterations. We choose to prioritize mitigating interference over maximizing balance for two reasons. First, interference is typically very difficult, or even impossible, to measure and therefore hard to correct for in the analysis phase. Imbalance, on the other hand, can be addressed with standard covariate adjustment methods. Second, for the sexual assault prevention intervention in our simulated experiment, we particularly want to avoid false null findings because this can keep an intervention that helps protect girls’ safety and well-being from being more widely implemented. Because interference often causes bias towards the null, we aim to prevent interference as much as possible. More generally, when deciding on a fitness function and restriction rule, study analysts may want to make methodological considerations, e.g. whether statistical corrections can be made during the analysis phase, and policy considerations, e.g. the consequences of a false negative or false positive finding.
4.2.5 Pre-registration
When the research team is satisfied with the choice of fitness function and restriction rule, they lock in the assignment mechanism $P_{IGR}$. They then pre-register the accepted allocations $\mathbf{z}^{j}$ in $\mathcal{Z}_{pool}$ that have $p_{\mathbf{z}_{j}}>0$ and the associated probabilities $p_{\mathbf{z}}$. For the purpose of guaranteeing valid statistical inference, we only need to pre-register these accepted allocations and probabilities. However, it can also be helpful to pre-register the fitness function $f$ and the restriction rule $r$ as an informative reference for the design of future trials. {leftrule}
Pre-registration for a Simulated Experiment
We pre-register the set of 500 accepted allocations, each with equal probability $\frac{1}{500}$ of being drawn. Additionally, we pre-register the fitness function $f^{(\texttt{F+M})}$ and threshold restriction rule $r$.
4.2.6 Randomization
Finally, the observed allocation $\mathbf{z}^{obs}\sim P_{IGR}$ is drawn and deployed. {leftrule}
Randomization in a Simulated Experiment
We sample one of the 500 accepted allocations uniformly at random.
5 Analyzing IGR-Designed Experiments
In this section, we examine bias properties of a difference-in-means estimate applied to an IGR-designed experiment. We first discuss bias under the usual no-interference assumption, then expand the discussion to settings where interference is present.
5.1 IGR Bias Properties
5.1.1 Unbiasedness Through the Mirror Property
@@
We show that IGR produces an unbiased difference-in-means estimate of the sample average treatment effect if there is no interference, if the IGR assignment mechanism satisfies the mirror property, and if in every accepted allocation, the number of treated and control individuals are equal. For a binary intervention, the mirror property is satisfied if $P(Z_{i}=0)=P(Z_{i}=1)$ for all $i=1,\dots,N$. Suppose we enumerated all possible allocations in the Enumeration step. If we specify fitness function $f$ and restriction rule $r$ to be symmetric, such that $(r\circ f)(\mathbf{X},\mathbf{z}))=(r\circ f)(\mathbf{X},\mathbf{1}-\mathbf{z}))$, the mirror property will be satisfied by $P_{IGR}$. However, because we may not be able to enumerate all possible allocations, satisfaction of the mirror property is not guaranteed. Instead, we can update $P_{IGR}$ by ensuring that, for every allocation that is accepted, its mirror allocation is also accepted. Let $\tilde{\mathcal{Z}}_{pool}=\{1-\mathbf{z}\mid\mathbf{z}\in\mathcal{Z}_{pool}\}$ denote the set of mirror allocations for each allocation in $\mathcal{Z}_{pool}$. The assignment mechanism $\tilde{P}_{IGR}$ satisfies the mirror property, where
@@
Proposition 1.
Assume that there is no interference and that the consistency property holds, so $Y_{i}=Y_{i}(1)Z_{i}+Y_{i}(0)(1-Z_{i}).$ If the assignment mechanism $\tilde{P}_{IGR}$ satisfies the mirror property and assigns an equal number of individuals to treatment and control in every allocation, then the difference-in-means estimate for the sample average treatment effect is unbiased.
Proof.
See Appendix F.1. ∎
5.1.2 Bias in the Presence of Interference
In an experiment with interference where the target estimand is the global average treatment effect, a design that satisfies the mirror property does not necessarily give an unbiased difference-in-means estimate because the interference itself introduces bias. To reduce interference bias, we may need to use a design that violates the mirror property. When the mirror property is violated, however, the individual’s treatment assignment and potential outcome may no longer be independent, resulting in possible bias from confounding. It may nevertheless may be beneficial to break the mirror property so long as doing so produces a sufficient reduction in interference bias.
In IGR, violation of the mirror property occurs when we choose an asymmetric fitness function and restriction rule pair to restrict the design. To understand why an asymmetric restriction can help to prevent interference, we consider a toy example. Suppose a subset of three or more individuals in the study are connected by a star-shaped network, such that there is a single central individual who is connected to every outer individual, and all outer individuals are only connected to this central individual. Suppose further that, for a unit assigned to the control-level, exposure to treatment occurs if at least two neighbors are treated. Then, in the allocation where the center individual is treated and the outer individuals are control, none of the control individuals are exposed. Conversely, in the mirror allocation where the center individual is control and the outer individuals are treated, the center control individual becomes exposed. The fitness function and restriction rule could be specified in such a way that, in order to rule out high-interference allocations, the latter allocation is rejected while the former is accepted. In Appendix F.2, we derive conditions under which asymmetric restriction in IGR improves overall bias for a linear outcomes model with an additive direct effect and an additive neighborhood exposure spillover effect.
Note that, even if the fitness function and restriction rule are asymmetric, we can force the IGR assignment mechanism to satisfy the mirror property by incorporating all mirror allocations as “accepted” allocations, regardless of whether they actually satisfy the criteria for acceptance. The analyst may consider enforcing the mirror property if the difference between the scores $f(\mathbf{X},\mathbf{z})$ and $f(\mathbf{X},\mathbf{1-z})$ is small for the $\mathbf{z}$’s in the candidate pool of allocations. We empirically investigate the implications of including and excluding mirror allocations in Section 6.
5.2 Randomization Inference for IGR-Designed Experiments